OpenStudy (anonymous):

What is the direct variation equation if y varies directly with x and y = –6 when x = –36? A.y=-6x B.y= 1/6x C.y=6x D.y=-1/6x

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@bibby @DanJS

OpenStudy (danjs):

Y varies directly with x. \[y ~\alpha ~x\] y = k*x Given (x,y) = (-36 , -6). Put those into the direct variation, and calculate the constant k.

OpenStudy (danjs):

A direct variation, y=k*x , is an equation for a line that goes through the origin, (0,0)

OpenStudy (danjs):

And in this case, also through the point (x,y) = (-36, -6).

OpenStudy (anonymous):

if im following you right and if im doing my math right it should be C?

OpenStudy (danjs):

\[-6 = k*-36\] \[k = \frac{ -6 }{ -36 } = \frac{ 1 }{ 6 }\]

OpenStudy (danjs):

\[y = \frac{ 1 }{ 6 }x\]

OpenStudy (anonymous):

Ah ha thanks so much ^.^

OpenStudy (danjs):

You just mixed up the x and y, that is what C is, if you plug them in backwards.

OpenStudy (anonymous):

yea thats what i figured but thanks i will tag you if i need more help

OpenStudy (danjs):

ok